# Deep Incentive Design with Differentiable Equilibrium Blocks

### 08 Mar 2026 v1

@@ 1 Introduction
We focus on correlated and coarse correlated equilibria as the lower-level constraints. These are natural choices for a problem setting where a central designer exists and can help correlate players’ strategies. Moreover, mathematically, these solution concepts allow us to select the unique maximum-entropy equilibrium from the convex equilibrium set, rendering this choice differentiable with respect to $\theta$.
@@
  • Experimentally, we show the promise of DID by tackling several well-studied and diverse problems from the literature, including multi-agent contract design [Holmstrom, 1982], machine scheduling [Heydenreich et al., 2007] and inverse equilibrium problems [Kuleshov and Schrijvers, 2015].
@@ 2.1 Gradient-Based MPEC
Several works have studied gradient-based approaches to solve MPECs. Generally these can be grouped into two approaches: those using implicit differentiation, relying on the implicit function theorem, and those using automatic differentiation, unrolling the computational graph of some converging dynamics. Both approaches rely on assuming a unique Nash equilibrium exists. Implicit differentiation has been used to converge to unique Stackelberg equilibria [Fiez et al., 2020, Wang et al., 2022]. Li et al. [2020] study (ID) for a single, fixed context and frame the lower-level equilibrium problem as a monotone variational inequality, which they solve with both implicit and automatic differentiation. Extensions to this work have included distributed algorithms [Grontas et al., 2024], hierarchy-free two-timescale methods [Liu et al., 2022, Li et al., 2023, Maheshwari et al., 2022], and using automatic differentiation to design congestion games [Li et al., 2024], mean-field games [Corecco et al., 2025], and social-dilemmas [Baumann et al., 2018, Balaguer et al., 2022].
@@ 2.2 Differentiable Economics
Conitzer and Sandholm [2002, 2004] were the first to propose automated mechanism design, i.e., the (computational) design of a tailored mechanism to the specific problem. More recently, Dütting et al. [2024] introduced differentiable economics, using neural networks to parameterize an auction mechanism. The network takes as input the bids and outputs the payments and allocations. Some constraints, such as payment non-negativity and allocation feasibility are hardcoded in the architecture, others such as strategyproofness are enforced via a Lagrangian penalty. Several works have since proposed improved training pipelines and architectures [Duan et al., 2022, Tacchetti et al., 2022, Ivanov et al., 2022, Rahme et al., 2021, Wang et al., 2024, Duan et al., 2023, Curry et al., 2023, 2024]. We refer to Curry et al. [2025] for a comprehensive survey of the field.
@@ 3.1 Game Theory and Equilibria
$\sigma^{*}$ is an $\varepsilon$-correlated equilibrium ($\varepsilon\text{-CE}$) [Aumann, 1974], if no player achieves more than an $\varepsilon$ gain in expected payoff by deviating to any action $a_{p}^{\prime}$, after having received a recommended action $a_{p}^{\prime\prime}$, drawn from $\sigma^{*}$. More formally, $\sigma^{*}\in\varepsilon\text{-CE}(G)$, if for all $p\in[1,N],a^{\prime}_{p}\in\mathcal{A}_{p},a^{\prime\prime}_{p}\neq a^{\prime}_{p}\in\mathcal{A}_{p}$:
@@
In case players can only choose to deviate before observing the recommended action, we call $\sigma^{*}$ an $\varepsilon$-coarse correlated equilibrium ($\varepsilon\text{-CCE}$) [Moulin and Vial, 1978]. Formally, $\sigma^{*}\in\varepsilon\text{-CCE}(G)$, if for all $p\in[1,N],a^{\prime}_{p}\in\mathcal{A}_{p}$:
@@
Both equilibrium notions require a common source of randomness to act as a coordination mechanism. This need not be an explicit signal, but could be a shared history of play. In fact, in a repeated game setup, no-regret learning algorithms, such as regret matching, use this history to provably converge to the set of (coarse) correlated equilibria, further motivating their use in this work [Hart and Mas-Colell, 2000].
A (C)CE is also a Nash equilibrium (NE), if it further holds that the joint can be written as the product of independent player marginals, i.e., if the players choose their actions independently. Note that $\text{NE}\subseteq\text{CE}\subseteq\text{CCE}$ and $\varepsilon_{1}\text{-(C)CE}\subseteq\varepsilon_{2}\text{-(C)CE}$ for $\varepsilon_{1}\leq\varepsilon_{2}$. As every normal-form game has a NE [Nash, 1951], it follows that none of the equilibrium sets is empty. Further note that the set of $\varepsilon$-(C)CEs always is a convex polytope, which is generally not true for NEs, with the exception of two-player constant-sum games when $\varepsilon=0$. Our DID framework requires the equilibrium set being a convex polytope. However, the approach works analogously for $\varepsilon$-CEs and $\varepsilon$-CCEs. [^1] DID could also be extended to implementing other equilibria, as long as their equilibrium set is a convex polytope, such as communication equilibria or Bayes-correlated equilibria. We therefore write $\varepsilon$-Eql(G)(G) to denote the convex equilibrium polytope of $G$ for statements that hold equivalently for both $\varepsilon$-CEs and $\varepsilon$-CCEs.
While the NE is often used in the literature, it is not suitable for our purposes. In particular, there are several obstacles to tackling (ID) with Nash constraints via gradient-based optimization. These include the computational hardness [Daskalakis et al., 2009], the multiplicity of NEs, and the fact that for general-sum games the set of NEs is not necessarily connected. In addition, as we vary $\theta$, certain connected components of the NE set may disappear and new ones appear [Kohlberg and Mertens, 1986].
@@ 5.1 Network Architecture
For two of the three experiments—the inverse equilibrium and machine scheduling problems—the mechanism generators take as input the initial payoffs and output perturbed payoffs of the same shape $N,|\mathcal{A}_{1}|,\dots,|\mathcal{A}_{N}|]$. [^6] In these cases, the context is equal to the initial payoffs. The details can be found in Sections 5.4 and 5.5 and Appendix A. These are of size $N\prod_{p=1}^{N}|\mathcal{A}_{p}|$, presenting a challenge for training on large games. Fortunately, the game-theoretic maps we are trying to learn have many equivariances. A function is equivariant to a permutation $\tau$ if it commutes with it, i.e., $f(\tau(x))=\tau(f(x)).$ In our case, any permutation of either the players or their action spaces in the input payoffs should also commute with the map $G(\theta;\omega)$, i.e., $G(\theta;\omega)$ is equivariant with respect to these permutations. Building a network architecture that respects these equivariances gives the network a strong inductive bias, as every input represents all possible $N!\prod_{p=1}^{N}(|\mathcal{A}_{p}|!)$ permutations of itself. As outlined below, an equivariant architecture also massively reduces the number of trainable parameters in the network, resulting in a more scalable training pipeline.
@@ 5.3 Multi-Agent Contract Design
Multi-Agent Contract Design was introduced by Holmstrom [1982]. Subsequent works focused on the complexity of implementing equilibria in multi-agent contracts under different assumptions such as linear contracts or binary vs. general action spaces [Babaioff and Winter, 2014, Dütting et al., 2023, 2025b, Cacciamani et al., 2024, Castiglioni et al., 2023]. Recently, Dütting et al. [2025a] introduced a model of multi-agent contract design where the lower-level problem is a (coarse) correlated equilibrium, similar to the $\varepsilon$-ME-Eql solution concept we use in this work. Despite the great economic interest in contract design, we are not aware of existing, implementable algorithms for the multi-agent case, underlining the promise of DID.
@@
Consider the contract design problem illustrated in Figure˜2 as an intuitive example. There are two players “Alice” and “Bob”, who are siblings and tasked to set up a Christmas tree. They each have two actions, they can either help or not help set up the tree and will incur costs depending on their exerted efforts. Their choices stochastically lead to one of three outcomes: (i) the tree is set up if they either both helped or with a $50\%$ chance if one helped, (ii) the tree is set up wrong and burns down with a $50\%$ chance if only one child helped, and (iii) no tree is set up if no one helps. The father prefers the tree being set up over not being set up over the tree burning down. Importantly, the father comes home and will not know who helped and who did not, i.e., he cannot observe the joint action $a$. He can only observe the tree, i.e., the outcome $o$, and promise to pay his children contracts contingent on the observed outcome.
@@ 5.4 Inverse Equilibrium Problems
Here we are interested in the following variant of the problem. Given a target equilibrium ${\sigma}^{\omega}$, which defines the context $\omega$, generate a game $G(\theta;\omega)$, such that its $\varepsilon$-ME-Eql minimizes the KL divergence to ${\sigma}^{\omega}$. Such problems that aim to compute game-theoretic equilibria, which are anchored to learned policies from (potentially human) expert behavior, are critical for many diverse ML applications, as they can be used, e.g., to discover near-optimal game play that is also interpretable and natural [Bakhtin et al., 2022, Jacob et al., 2022, Lerer and Peysakhovich, 2019, McIlroy-Young et al., 2020, Munos et al., 2024]. This problem can be formulated as an instance of ($\varepsilon$-ME-ID) as follows:
@@
We evaluate on 8 samples for each game with shape in $\{2{\times}2,\dots,16{\times}16\}$ and report the average over all contexts. Due to the increased dimensionality of the search space, we polish using 4,000 function evaluations. The results can be found in Table˜1(b). The average KL divergence of our approach significantly outperforms the naive benchmark of outputting a constant game with the uniform joint as $\varepsilon$-ME-Eql. As before, there is some room for improvement if we polish the solution using a local optimizer. Unsurprisingly, we also see reduced performance when evaluating under ECOS, highlighting the approximate nature of the DEB. Note that for $2{\times}2$ games, we can identify the joint strategy space with $\Delta^{3}$ and thus plot our results on the simplex. Figure˜3 illustrates two sampled targets $\sigma^{w}$, the $\varepsilon$-Eql polytope of the generated game $G(\theta,\omega)$ and the resulting $\varepsilon$-ME-Eql $\sigma^{*}$, which nicely matches the target.
@@ 5.5 Machine Scheduling
We consider an incentive designer that wants to minimize the maximum makespan of the machines given by $\max_{j}\sum_{p:a_{p}=j}t^{\omega}_{p,j}$. We assume the designer can levy taxes $s_{p}(a;\theta,\omega)\geq 0$ on the agents, i.e., positive transfers from the agents to the designer, but wants to keep these minimal. The designer thus induces a game $G(\theta,\omega)$ with payoffs $G_{p}(\theta;\omega)=u_{p}(a;\omega)-s_{p}(a;\theta)$. Formulating the machine scheduling problem as an instance of ($\varepsilon$-ME-ID) yields
@@ 6 Conclusion
Section˜5 illustrated both the success of our approach on three diverse tasks and, more importantly, how easily it can be adapted to different problems. Indeed, with one single DEB (per desired solution concept), one fundamental type of equivariant layer, and a unified training pipeline with many fixed hyperparameters, we were able to train mechanism generators for three different problems, which themselves generalized across many game sizes and contexts. Given this generalizability, we anticipate that DID will be useful across the multi-agent community in tackling a wide array of problems with equilibrium constraints.
@@ A.1 Architectures and Code
Our networks are built to respect the equivariances of the underlying game-theoretic functions. A function $f$ is said to be equivariant if applying a permutation $\tau$ to its input is equivalent to applying the same permutation to the output, i.e.,
@@
The game-theoretic functions we are trying to learn have several such equivariances, e.g., if we permute the action space of a player $p$ in a game $G$ to yield a game $G_{\tau}$ and do the same for $G$’s equilibrium $\sigma^{*}$ to yield $\sigma^{*}_{\tau}$ then $\sigma^{*}_{\tau}\in\text{Eql}(G_{\tau})$ and we say the equilibrium map is equivariant with respect to the permutation $\tau$.
We briefly explain the fundamental architecture used for all layers in our network. All our data has a channel dimension and a varying amount of equivariant dimensions (such as players, strategies and outcomes). For a given input at channel $c$, $(c,\dots)$ we apply a set of pooling functions $\phi_{i}$ which preserve some corresponding equivariances $\tau_{i}$ of the equivariant dimensions, i.e.,
@@
The output from the pooling functions is then concatenated along the channel dimension and passed through a linear transform, whose weights are shared across the equivariant dimensions, i.e., only depend on the channel dimension. This ensures the equivariance is preserved and gives the network the inductive bias. Afterwards we apply a (nonlinear) activation function. In general we can thus recursively define the output $g_{l+1}$ of the $l+1$ layer as
@@ A.2.1 Architecture
In our experiments we use three hidden layers with 64 nodes each, i.e., $C_{l}=64$. Note that after the pooling we have $C_{l}I=512$ channels, which then get mapped back to $C_{l+1}=64$ channels by the linear layer. The exception is the first layer, which only has $2$ input channels and the final layer, whose output has channel dimension $1$. Moreover, for the final layer we do not apply a GELU activation but the identity to ensure the network can generate both positive and negative payoffs.

### 12 Mar 2026 v2

@@ 1 Introduction
We focus on correlated and coarse correlated equilibria as the lower-level constraints. It is a natural choice for a problem setting where a central designer exists and can help correlate players’ strategies. Moreover, mathematically, it allows us to select the unique maximum-entropy equilibrium from the convex equilibrium set, rendering this choice differentiable with respect to $\theta$.
@@
  • Experimentally, we show the promise of DID by tackling several well-studied and diverse problems from the literature, including multi-agent contract design Holmstrom [1982], machine scheduling Heydenreich et al. [2007] and inverse equilibrium problems Kuleshov and Schrijvers [2015].
@@ 2.1 Gradient-Based MPEC
Several works have studied gradient-based approaches to solve MPECs. Generally these can be grouped into two approaches: those using implicit differentiation, relying on the implicit function theorem, and those using automatic differentiation, unrolling the computational graph of some converging dynamics. Both approaches rely on assuming a unique Nash equilibrium exists. Implicit differentiation has been used to converge to unique Stackelberg equilibria Fiez et al. [2020], Wang et al. [2022]. Li et al. [2020] study (ID) for a single, fixed context and frame the lower-level equilibrium problem as a monotone variational inequality, which they solve with both implicit and automatic differentiation. Extensions to this work have included distributed algorithms Grontas et al. [2024], hierarchy-free two-timescale methods Liu et al. [2022], Li et al. [2023], Maheshwari et al. [2022], and using automatic differentiation to design congestion games Li et al. [2024], mean-field games Corecco et al. [2025], and social-dilemmas Baumann et al. [2018], Balaguer et al. [2022].
@@ 2.2 Differentiable Economics
Conitzer and Sandholm [2002, 2004] were the first to propose automated mechanism design, i.e. the (computational) design of a tailored mechanism to the specific problem. More recently, Dütting et al. [2024] introduced differentiable economics, using neural networks to parameterize an auction mechanism. The network takes as input the bids and outputs the payments and allocations. Some constraints, such as payment non-negativity and allocation feasibility are hardcoded in the architecture, others such as strategyproofness are enforced via a Lagrangian penalty. Several works have since proposed improved training pipelines and architectures [Duan et al., 2022, Tacchetti et al., 2022, Ivanov et al., 2022, Rahme et al., 2021, Wang et al., 2024, Duan et al., 2023, Curry et al., 2023, 2024]. We refer to Curry et al. [2025] for a comprehensive survey of the field.
@@ 3.1 Game Theory and Equilibria
$\sigma^{*}$ is an $\varepsilon$-correlated equilibrium ($\varepsilon\text{-CE}$) Aumann [1974], if no player achieves more than an $\varepsilon$ gain in expected payoff by deviating to any action $a_{p}^{\prime}$, after having received a recommended action $a_{p}^{\prime\prime}$, drawn from $\sigma^{*}$. More formally, $\sigma^{*}\in\varepsilon\text{-CE}(G)$, if for all $p\in[1,N],a^{\prime}_{p}\in\mathcal{A}_{p},a^{\prime\prime}_{p}\neq a^{\prime}_{p}\in\mathcal{A}_{p}$:
@@
In case players can only choose to deviate before observing the recommended action, we call $\sigma^{*}$ an $\varepsilon$-coarse correlated equilibrium ($\varepsilon\text{-CCE}$) Moulin and Vial [1978]. Formally, $\sigma^{*}\in\varepsilon\text{-CCE}(G)$, if for all $p\in[1,N],a^{\prime}_{p}\in\mathcal{A}_{p}$:
@@
Both equilibrium notions require a common source of randomness to act as a coordination mechanism. This need not be an explicit signal, but could be a shared history of play. In fact, in a repeated game setup, no-regret learning algorithms, such as regret matching, use this history to provably converge to the set of (coarse) correlated equilibria, further motivating their use in this work Hart and Mas-Colell [2000].
A (C)CE is also a Nash equilibrium (NE), if it further holds that the joint can be written as the product of independent player marginals, i.e. if the players choose their actions independently. Note that $\text{NE}\subseteq\text{CE}\subseteq\text{CCE}$ and $\varepsilon_{1}\text{-(C)CE}\subseteq\varepsilon_{2}\text{-(C)CE}$ for $\varepsilon_{1}\leq\varepsilon_{2}$. As every normal-form game has a NE Nash [1951], it follows that none of the equilibrium sets is empty. Further note that the set of $\varepsilon$-(C)CEs always is a convex polytope, which is generally not true for NEs, with the exception of two-player constant-sum games when $\varepsilon=0$. Our DID framework requires the equilibrium set being a convex polytope. However, the approach works analogously for $\varepsilon$-CEs and $\varepsilon$-CCEs. [^1] DID could also be extended to implementing other equilibria, as long as their equilibrium set is a convex polytope, such as communication equilibria or Bayes-correlated equilibria. We therefore write $\varepsilon$-Eql(G)(G) to denote the convex equilibrium polytope of $G$ for statements that hold equivalently for both $\varepsilon$-CEs and $\varepsilon$-CCEs.
While the NE is often used in the literature, it is not suitable for our purposes. In particular, there are several obstacles to tackling (ID) with Nash constraints via gradient-based optimization. These include the computational hardness Daskalakis et al. [2009], the multiplicity of NEs and that for general-sum games the set of NEs is not necessarily connected. To add to that, as we vary $\theta$ certain connected components of the NE set may disappear and new ones appear Kohlberg and Mertens [1986].
@@ 5.1 Network Architecture
For two of the three experiments–the inverse equilibrium and machine scheduling problems–the mechanism generators take as input the initial payoffs and output perturbed payoffs of the same shape $N,|\mathcal{A}_{1}|,\dots,|\mathcal{A}_{N}|]$. [^6] In these cases, the context is equal to the initial payoffs. The details can be found in Sections 5.4 and 5.5 and Appendix A. These are of size $N\prod_{p=1}^{N}|\mathcal{A}_{p}|$, presenting a challenge for training on large games. Fortunately, the game-theoretic maps we are trying to learn have many equivariances. A function is equivariant to a permutation $\tau$ if it commutes with it, i.e. $f(\tau(x))=\tau(f(x)).$ In our case, any permutation of either the players or their action spaces in the input payoffs should also commute with the map $G(\theta;\omega)$, i.e. $G(\theta;\omega)$ is equivariant with respect to these permutations. Building a network architecture that respects these equivariances gives the network a strong inductive bias, as every input represents all possible $N!\prod_{p=1}^{N}(|\mathcal{A}_{p}|!)$ permutations of itself. As outlined below, an equivariant architecture also massively reduces the number of trainable parameters in the network, resulting in a more scalable training pipeline.
@@ 5.3 Multi-Agent Contract Design
Multi-Agent Contract Design was introduced by Holmstrom [1982]. Consecutive works focused on the complexity of implementing equilibria in multi-agent contracts under different assumptions such as linear contracts or binary vs. general action spaces Babaioff and Winter [2014], Dütting et al. [2023, 2025b], Cacciamani et al. [2024], Castiglioni et al. [2023]. Recently, Dütting et al. [2025a] introduced a model of multi-agent contract design where the lower-level problem is a (coarse) correlated equilibrium, similar to the $\varepsilon$-ME-Eql solution concept we use in this work. Despite the great economic interest in contract design, we are not aware of existing, implementable algorithms for the multi-agent case, underlining the promise of DID.
@@
Consider the contract design problem illustrated in Figure˜2 as an intuitive example. There are two players “Alice” and “Bob”, who are siblings and tasked to set up a Christmas tree. They each have two actions, they can either help or not help set up the tree and will incur costs depending on their exerted efforts. Their choices stochastically lead to one of three outcomes: (i) the tree is set up if they either both helped or with a $50\%$ chance if one helped, (ii) the tree is set up wrong and burns down with a $50\%$ chance if only one child helped, and (iii) no tree is set up if no one helps. The father prefers the tree being set up over not being set up over the tree burning down. Importantly the father comes home and will not know who helped and who did not, i.e. he cannot observe the joint action $a$. He can only observe the tree, i.e. the outcome $o$ and promise to pay his children contracts contingent on the observed outcome.
@@ 5.4 Inverse Equilibrium Problems
Here we are interested in the following variant of the problem. Given a target equilibrium ${\sigma}^{\omega}$, which defines the context $\omega$, generate a game $G(\theta;\omega)$, such that its $\varepsilon$-ME-Eql minimizes the KL divergence to ${\sigma}^{\omega}$. Such problems that aim to compute game-theoretic equilibria, which are anchored to learned policies from (potentially human) expert behavior are critical for many diverse ML applications as e.g. they can be used to discover near-optimal game play that is also interpretable and natural [Bakhtin et al., 2022, Jacob et al., 2022, Lerer and Peysakhovich, 2019, McIlroy-Young et al., 2020, Munos et al., 2024]. This problem can be formulated as an instance of ($\varepsilon$-ME-ID) as follows:
@@
We evaluate on 8 samples for each game with shape in $\{2{\times}2,...,16{\times}16\}$ and report the average over all contexts. Due to the increased dimensionality of the search space, we polish using 4,000 function evaluations. The results can be found in Table˜1(b). The average KL divergence of our approach significantly outperforms the naive benchmark of outputting a constant game with the uniform joint as $\varepsilon$-ME-Eql. As before, there is some room for improvement if we polish the solution using a local optimizer. Unsurprisingly, we also see reduced performance when evaluating under ECOS, highlighting the approximate nature of the DEB. Note that for $2{\times}2$ games, we can identify the joint strategy space with $\Delta^{3}$ and thus plot our results on the simplex. Figure˜3 illustrates two sampled targets $\sigma^{w}$, the $\varepsilon$-Eql polytope of the generated game $G(\theta,\omega)$ and the resulting $\varepsilon$-ME-Eql $\sigma^{*}$, which nicely matches the target.
@@ 5.5 Machine Scheduling
We consider an incentive designer that wants to minimize the maximum makespan of the machines given by $\max_{j}\sum_{p:a_{p}=j}t^{\omega}_{p,j}$. We assume the designer can levy taxes $s_{p}(a;\theta,\omega)\geq 0$ on the agents, i.e. positive transfers from the agents to the designer, but wants to keep these minimal. The designer thus induces a game $G(\theta,\omega)$ with payoffs $G_{p}(\theta;\omega)=u_{p}(a;\omega)-s_{p}(a;\theta)$. Formulating the machine scheduling problem as an instance of ($\varepsilon$-ME-ID) yields
@@ 6 Conclusion
Section˜5 illustrated both the success of our approach on three diverse tasks and more importantly how easily it can be adapted to different problems. Indeed, with one single DEB (per desired solution concept), one fundamental type of equivariant layer, and a unified training pipeline with many fixed hyperparameters, we were able to train mechanism generators for three different problems, which themselves generalized across many game sizes and contexts. Given this generalizability, we anticipate that DID will be useful across the multi-agent community in tackling a wide array of problems with equilibrium constraints.
@@
Acknowledgements
We would like to thank Ian Gemp, Siqi Liu, Paul Dütting, and Marc Lanctot for their constructive feedback and the helpful discussions during the writing of this paper.
@@ A.1 Architectures and Code
Our networks are built to respect the equivariances of the underlying game-theoretic functions. A function $f$ is said to be equivariant if applying a permutation $\tau$ to its input is equivalent to applying the same permutation to the output, i.e.
@@
The game-theoretic functions we are trying to learn have several such equivariances, e.g. if we permute the action space of a player $p$ in a game $G$ to yield a game $G_{\tau}$ and do the same for $G$’s equilibrium $\sigma^{*}$ to yield $\sigma^{*}_{\tau}$ then $\sigma^{*}_{\tau}\in\text{Eql}(G_{\tau})$ and we say the equilibrium map is equivariant with respect to the permutation $\tau$.
We briefly explain the fundamental architecture used for all layers in our network. All our data has a channel dimension and a varying amount of equivariant dimensions (such as players, strategies and outcomes). For a given input at channel $c$, $(c,\dots)$ we apply a set of pooling functions $\phi_{i}$ which preserve some corresponding equivariances $\tau_{i}$ of the equivariant dimensions, i.e.
@@
The output from the pooling functions is then concatenated along the channel dimension and passed through a linear transform, whose weights are shared across the equivariant dimensions, i.e. only depend on the channel dimension. This ensures the equivariance is preserved and gives the network the inductive bias. Afterwards we apply a (nonlinear) activation function. In general we can thus recursively define the output $g_{l+1}$ of the $l+1$ layer as
@@ A.2.1 Architecture
In our experiments we use three hidden layers with 64 nodes each, i.e. $C_{l}=64$. Note that after the pooling we have $C_{l}I=512$ channels, which then get mapped back to $C_{l+1}=64$ channels by the linear layer. The exception is the first layer, which only has $2$ input channels and the final layer, whose output has channel dimension $1$. Moreover, for the final layer we do not apply a GELU activation but the identity to ensure the network can generate both positive and negative payoffs.